A share is yielding constant dividends for 5 years and is going to be sold for 400 € at the beginning of the 6th year. The annual dividend is equal to 50 €. The issuing company is financed both by equity and debt and has a WACC of 8%. The dividend yield for 5 years and the financial value of the share, are: "| with steps "


a. 180 € and 139 €

 b. 250 € and 472 €

c. 170 € and 200 €

d. 560 € and 325 €

e. 240 € and 235 €


1
Expert's answer
2022-02-10T13:37:01-0500

"dividend yield=5\times250"

"financial value of the share=\frac{50}{(1+0.08)^1}+\frac{50}{(1+0.08)^2}+\frac{50}{(1+0.08)^3}+\frac{50}{(1+0.08)^4}+\frac{50}{(1+0.08)^5}+\frac{400}{(1+0.08)^1}=471.86=472"


b.